OpenStudy (anonymous):

Can we solve this? 1 ∫ (x ln x )dx 0

OpenStudy (anonymous):

\[\int\limits_{1}^{0} x lnx dx\]

OpenStudy (anonymous):

parts is my guess

OpenStudy (anonymous):

improper integral, so it may not exist, but you can find the anti derivative by parts and then check

OpenStudy (anonymous):

ooh that was wrong sorry

OpenStudy (anonymous):

we can solve it

OpenStudy (anonymous):

\[\int udv = uv-\int vdu\] with \(u=\ln(x), du =\frac{dx}{x}, dv = x, v = \frac{x^2}{2}\) should do it

OpenStudy (anonymous):

"anti derivative" will be \[\frac{x^2}{2}\ln(x)-\frac{x^2}{4}\] i think

OpenStudy (anonymous):

now you are going to need to find that \(\lim_{x\to 0}\frac{x^2}{2}\ln(x)=0\)

OpenStudy (anonymous):

I know that we can solve it by parts, but due to it's definite integral 0 and 1, so i think ln(0) is not possible.

OpenStudy (anonymous):

when i tried this in wolfrom mathematica, it was giving answer of -1/4.

OpenStudy (anonymous):

like satellite73 said, that's the answer... what you have to do is, find the limit \[\lim_{x \rightarrow 0} \frac{ x^{2} }{ 2 } \ln x\]... I think that's done if you know logarithmic series...

OpenStudy (anonymous):

and you use L'Hopital's rule then.. try plugging in the logarithmic infinte series for log x...hope this hint helps..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

hey, can u say, how to apply L'Hospital's Rule in finding Log[0]?